Answer the following conceptual questions and practice about potential and kinetic energy.


1. You have two balls, a tennis ball and a baseball with twice the mass of the tennis ball. If you want to throw both balls with the same kinetic energy, how much faster should the tennis ball be compared to the baseball?


1
Expert's answer
2021-11-22T12:37:58-0500

"mv^2/2=(2m)u^2/2\to v^2=2u^2 \to v=u\sqrt{2}"


"v" is the speed of a tennis ball

"u" is the speed of a baseball ball